Adigirl29:

Show all work to solve the equation for x. If a solution is extraneous, be sure to identify it in your final answer. square root of the quantity x plus 7 end quantity minus 1 equals x

Adigirl29:

nope-

Adigirl29:

um how would I do that?

Adigirl29:

oh so then minus 1 would be -1

Adigirl29:

what about the quantity though?

Adigirl29:

what about the quantity tho

Adigirl29:

what do I put for that?

Adigirl29:

ohhhhh that makes more sense

Adigirl29:

do we solve what's in the parenthesis first?

Adigirl29:

I think so?

Adigirl29:

oh so we had to remove them

Adigirl29:

so would I need a calculator?

Adigirl29:

so would we add 10 plus 25?

Adigirl29:

yeah

Adigirl29:

ohhh now I get it

Adigirl29:

wait so would I have to write the long version too?

Adigirl29:

do I just write down what the link steps are?

Adigirl29:

oh alright so I would write down solving it with the quadratic formula too

Adigirl29:

oh

Adigirl29:

ill just write down the steps it shows

Adigirl29:

test how?

Adigirl29:

oh so I plug it in

Adigirl29:

what do I do once I plug it in?

snowflake0531:

@astro wrote:
I'll write it for you... \[\sqrt{(x+7-1)=x}\]
respectfully disagree it would be \[\sqrt{x+7}-1=x\]

Adigirl29:

um

snowflake0531:

It's two entirely different questions with the 1 in and out

snowflake0531:

it's not correct, if you had -1 under the square root, it would be square root x + 6, if you had in outside however... you would first need to add 1 to both sides, THEN square

Adigirl29:

im very confused-

snowflake0531:

add 1 to both sides \[\sqrt{x+7} = x+1 \\x+7 = (x+1)^2\]

Adigirl29:

wait so who's right then

Timmyspu:

If I am looking and solving this correctly as snow goes. I would have to say snow is correct.

Adigirl29:

so the other guy was wrong? yikes

Adigirl29:

what would be the correct way to solve this then cus now im confused

snowflake0531:

"square root of the quantity x plus 7 end quantity minus 1 equals x" think of the quantity and end quantity as parenthesis square root of the quantity x plus seven, end quantity is \(\sqrt{x+7}\) then that minus 1, equals to x

snowflake0531:

then you get the equation \[\sqrt{x+7}-1=x\]

Adigirl29:

so I just type what you wrote

snowflake0531:

that would be the equation do you know how to solve for x?

Adigirl29:

no

snowflake0531:

can you add 1 to both sides

Adigirl29:

so for 7 and 1?

snowflake0531:

can you add the left hand expression by 1, and the right by 1

Adigirl29:

oh so the 2 x's basically?

snowflake0531:

do you understand how I get from \(\sqrt{x+7}-1=x\) to \(\sqrt{x+7}=x+1\)

Adigirl29:

we just add a one for each side

snowflake0531:

correct

snowflake0531:

now can you write out the equation you would get if you squared both sides?

Adigirl29:

would it be the same just fully squared? Orr

snowflake0531:

?

snowflake0531:

just write out what you think

Adigirl29:

I think it would look the same

snowflake0531:

\(\sqrt{x+7}=x+1\) square both sides \((\sqrt{x+7})^2 = (x+1)^2\) correct?

Adigirl29:

im running out of time could you maybe just write out the steps for me please

snowflake0531:

\[x+7 = (x+1)^2\] FOIL method: \((a+b)^2 = a^2 + 2ab+b^2\) \[x+7=x^2 + 2x+1 \\ x^2 +x -6=0\\(x+3)(x-2)=0\]

snowflake0531:

I factored it for you, with this you will get two solutions

snowflake0531:

one is extraneous the other is the actual answer one of them is extraneous because if you put it back into the original question, it won't go through the square root. so basically, hint the positive one is the answer, the negative is the extraneous solution

Adigirl29:

alright thanks